Session: Jun 2009 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Urumqi (Hua Zhang at Sat Jun 6 00:08:11 2009) Success :

Medicina (Franco Fiocchi at Sat Jun 6 00:58:23 2009) Minor failures : RECO
Many "statisctics errors" reported form the Mk5.

Onsala (Michael Lindqvist at Sat Jun 6 06:01:48 2009) Success :
Apart from RFI and some rain. No known problems.

Westerbork (Geert Kuper at Sun Jun 7 18:23:32 2009) Success : LATE
Wb used in phased array mode. Source still byond ha limits at start. Tracking at 11:29:30 UT. RFI, further no known problems

Torun (Kaz Borkowski at Mon Jun 8 07:28:57 2009) Success : RFI
No known problems besides RFI.

Effelsberg (Uwe Bach at Mon Jun 8 08:20:53 2009) Success :

Jodrell_Bank (Alastair Gunn at Tue Jun 23 13:55:54 2009) Minor failures : ANT
The Lovell Telescope was used (Jb1). The Lovell AND Mk2 Telescopes were used. Lovell Telescope taken out of service at 11:23 UT due to discovery of tyre crack. Mk2 was substituted and schedule restarted at 12:34 UT. No other known problems.

Noto (P. Cassaro at Thu Jun 25 16:02:21 2009) Success :
No known problems